OpenStudy (anonymous):

find what c means - 27=3c-3(6-2c)

HanAkoSolo (jamierox4ev3r):

ah you want to find the value of c? first, lets simplify this equation together! That is what you are asking, isn't it...

OpenStudy (anonymous):

Yes Thanks!!!! Finally I get it Okay she said we need to simplify then use inverse then prove it and then do our properties

HanAkoSolo (jamierox4ev3r):

right, you simplify by distributing, add the like terms together, then you solve for c just like in a regular equation! :)

OpenStudy (anonymous):

okay so I don't think there is anything we can simplify

OpenStudy (anonymous):

Ok I guess that's correct thank you...... :)

OpenStudy (petiteme):

- 27=3c-3(6-2c) First, distribute -3 to the parenthesis so -27 = 3c-18+6c add like terms -27 = 9c - 18 To get c, transpose -18 to the other side -27 + 18 = 9c -9 = 9c divide both sides by 9 so c=-1

OpenStudy (anonymous):

wait that's not right but thank you I messed up and put -27 and it is just 27

HanAkoSolo (jamierox4ev3r):

thanks for explaining @petiteme , but I was hoping to get @haburke to figure it out by him/her self lol well that's okay, glad you get it now :)

OpenStudy (anonymous):

No it wasn't right Jaimerox4ev3r I messed up and put - 27

OpenStudy (petiteme):

Sorry :) @haburke if its not -27, just follow the steps :)

OpenStudy (anonymous):

ok thanks to you both

HanAkoSolo (jamierox4ev3r):

and plug the variable back into the equation, that should help you a lot! :) This will also help prove whether you are correct or not

OpenStudy (anonymous):

okay now I have another one but I got the answer and it works in the calculator but not when I try to prove it

OpenStudy (anonymous):

3+4(z+5)=3 I got z= 2

OpenStudy (petiteme):

distribute first 4 to the (z + 5)

OpenStudy (petiteme):

substitute z by 2 :)

HanAkoSolo (jamierox4ev3r):

lets double check by plugging in the values. 3+4(2+5)=3 3+8+20 does not equal 3....so that means taht 2 is not a correct answer. Let me walk you through a process similar to the one @petiteme showed earlier

HanAkoSolo (jamierox4ev3r):

So first, use distribution 3+4(z+5)=3 3+4z+20=3 Then add like terms 23+4z=3 Subtact 23 from both sides 23+4z=3 -23 -23 --------- 4z=-20 divide by 4 4z=-20 /4 /4 z=-5 ^ and that right there should be your answer :)

OpenStudy (anonymous):

ok see i did this : 3=4(z+15)=3 3+4z+20=31 4z+23=31 4z+23-23=31-23 4z=8 4z/4=8/4 z=2

HanAkoSolo (jamierox4ev3r):

how did you get 31 is my question? also, why do you have 15 in your top line inside the parentheses, i thought it was 5!. All of what you did seems correct, just not sure how you got 31.

OpenStudy (anonymous):

Okay the original problem was 3+4(z+5)=40

OpenStudy (anonymous):

crud no 31

OpenStudy (anonymous):

the original problem is 3+4(z+5)=31

HanAkoSolo (jamierox4ev3r):

wait so you wrote the problem wrong? O.o if so, rewrite it correctly please lol and then I can try to double check and solve it again :P if need be

OpenStudy (anonymous):

Okay I just cant type well today : 3+4(z+15)=31 3+4z+20=31 4Z+23=31 -23 -23 4Z= 8 4Z/4=8/4 Z=2

OpenStudy (anonymous):

??

OpenStudy (petiteme):

is it really 15 or 5?

OpenStudy (anonymous):

its 15 i just messed up sorry

OpenStudy (petiteme):

its okay :) if its 15, check your calculations. while distributing 4, 4(15) is not 20 :)